# Difference of surface effects in the inclusive and reactions

### v1

@@ IV SUMMARY
We have clarified the difference of the surface effects between the inclusive and reactions and the cause of the difference. The energy spectra calculated by SCDW with LFG and FG were compared to the experimental data of the and reactions on several combinations of targets and incident energies. By comparing the energy spectra with LFG and FG, we have found that the surface effect is necessary to reproduce the experimental data for both and reactions.
We have investigated in more details the difference of the surface effects in the two reactions by comparing the radial distributions of the energy spectra. The radial peak positions for each energy transfer shift to the outer region of the nucleus as decrease and the tendencies were almost the same for the two reactions. However, the opposite trend was observed between the two reactions in terms of the peak height. As a result, it has been found that the surface effect becomes larger for the reaction than the reaction. Moreover, we compared the radial distributions ignoring the absorption by the nucleus. The results show almost identical radial distributions and it has been found that the cause of the stronger surface effect in the reaction is the stronger absorption of deuteron than proton.
We have applied the above findings on surface effects in reaction to the improvement of the phenomenological model by Kalbach to calculate continuum deuteron inelastic scattering. We have newly introduced a surface effect into the model. For the finite well depth , which is the adjustable parameter associated with the surface effect, the optimized value for the proton-induced reactions was insufficient to reproduce the experimental data of the reaction. When the value of was set much smaller, as suggested by the analysis with SCDW, the calculated values reproduced the experimental data well over a wide emission energy range. This result has indicated that parameter optimization of a phenomenological model has been justified by a microscopic reaction model.
As one of the future works, it will be necessary to extend the present one-step SCDW model to describe the multi-step processes. With this extension, the SCDW model could guide the improvement of phenomenological models not only with respect to surface effects, which are prominent in the small region, but also with respect to wider energy spectra and angular distributions of outgoing deuterons. On another front, it is of interest to perform a similar analysis for other particles. -particle is a good candidate since it is well known to undergo strong absorption and is of large importance in the application fields.
@@ Acknowledgements.
- Kalbach [2000] C. Kalbach, Phys. Rev. C 62, 044608 (2000).

### v2

@@ IV SUMMARY
We have clarified the difference in the peripherality of the inclusive and reactions and the cause of the difference. The energy spectra calculated by SCDW with LFG and FG were compared to the experimental data of the and reactions on several combinations of targets and incident energies. By comparing the energy spectra with LFG and FG, we have found that the consideration of the nuclear surface is necessary to reproduce the experimental data for both and reactions.
We have investigated the difference in the peripherality of the and reactions by comparing the radial distributions of the energy spectra. The radial peak positions for each energy transfer shifts to the outer region of the nucleus as decrease and the tendencies were almost the same for the two reactions. However, the opposite trend was observed between the two reactions in terms of the peak height. As a result, it has been found that the peripherality get stronger for the reaction than the reaction. Moreover, we compared the radial distributions ignoring the absorption by the nucleus. The results show almost identical radial distributions and it has been found that the cause of the stronger peripherality of the reaction is the stronger absorption of deuteron than proton.
We have applied the above findings on the peripherality of reaction to the improvement of the phenomenological model by Kalbach to calculate continuum deuteron inelastic scattering. For the finite well depth , which is the adjustable parameter associated with the surface correction, the optimized value for the proton-induced reactions was insufficient to reproduce the experimental data of the reaction. When the value of was set much smaller, as suggested by the analysis with SCDW model, the calculated values reproduced the experimental data well over a wide emission energy range. This result has indicated that parameter optimization of the phenomenological reaction model has been justified by the SCDW model.
As one of the future works, it will be necessary to extend the present one-step SCDW model to describe the multi-step processes. With this extension, the SCDW model could guide the improvement of phenomenological models not only with respect to surface corrections, which are prominent in the small region, but also with respect to wider energy spectra and angular distributions of outgoing deuterons. On another front, it is of interest to perform a similar analysis for other particles. Ref. [14] was the first example to analyze the peripherality of the reaction in detail using the SCDW model. In this study, we presented that a similar analysis is also valid for reactions induced by deuteron. Future similar analyses for various composite particles will improve our understanding of the peripherality of nuclear reactions. -particle is a good candidate since it is well known to undergo strong absorption and is of large importance in the application fields.
@@ Acknowledgements.
- Griffin [1966] J. J. Griffin, Phys. Rev. Lett. 17, 478 (1966).
- Běták and Dobeš [1976] E. Běták and J. Dobeš, Z Physik A 279, 319–324 (1976).
- Avrigeanu et al. [1996] M. Avrigeanu, A. Harangozo, V. Avrigeanu, and A. N. Antonov, Phys. Rev. C 54, 2538 (1996).
- Avrigeanu et al. [1997] M. Avrigeanu, A. Harangozo, V. Avrigeanu, and A. N. Antonov, Phys. Rev. C 56, 1633 (1997).
@@
- Kalbach [1995] C. Kalbach, J. Phys. G Nucl. Part. Phys. 21, 1519 (1995).